# Rank-One Modified Value Iteration

### v1

@@ Abstract.
The authors are with Delft Center for Systems and Control, Delft University of Technology, Delft, The Netherlands. Email: {a.sharifikolarijani,t.ok,p.mohajerinesfahani,m.a.sharifikolarijani}@tudelft.nl.
This work was supported by the Horizon Europe Pathfinder Open project RELIEVE-101099481 and by the European Research Council (ERC) project TRUST-949796.
@@ 1. Introduction
Contribution. In this paper, we propose a novel algorithm that modifies the VI algorithm by incorporating a computationally efficient PI-type update rule. To this end, we consider the update rule (1) with a matrix gain of the form $\bm{G}_{k}=(\bm{I}-\widetilde{\bm{P}}_{k})^{-1}$, where $\widetilde{\bm{P}}_{k}$ is a rank-one approximation of $\bm{P}_{k}$. To be precise, we consider the approximation $\widetilde{\bm{P}}_{k}=\bm{1}\bm{d}_{k}^{\top}$, where $\bm{d}_{k}=\bm{P}_{k}^{\top}\bm{d}_{k}$ is the stationary distribution of $\bm{P}_{k}$ (assuming it exists) and $\bm{1}$ is the all-one vector. The proposed algorithm then uses the power method for approximating $\bm{d}_{k}$ iteratively using the true matrix $\bm{P}_{k}$ in the planning problem and its sampled version in the learning problem. In particular,
@@
    we compare the proposed R1-VI and R1-QL algorithms with the state-of-the-art algorithms for solving planning and learning problems of MDPs and show the empirically faster convergence of the proposed algorithms compared the ones with the same per-iteration computational complexity (i.e., the first-order algorithms and their accelerated versions).
@@ 2. Optimal control of MDPs
Consider a finite MDP $(\mathcal{S},\mathcal{A},P,c,\gamma)$. Here, $\mathcal{S}:=\{1,2,\ldots,n\}$ and $\mathcal{A}:=\{1,2,\ldots,m\}$ are the state and action spaces, respectively. The transition kernel $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the conditional probability $P(s^{+}|s,a)$ of the transition to state $s^{+}$ given the current state-action pair $(s,a)$. The function $\bm{c}\in\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|}=\mathbb{R}^{nm}$, bounded from below, represents the stage cost $\bm{c}(s,a)$ of taking the control action $a$ while the system is in state $s$. And, $\gamma\in(0,1)$ is the discount factor which can be seen as a trade-off parameter between short- and long-term costs. A control policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ is a mapping from states to actions. Fix policy $\pi$. For the corresponding Markov chain under the policy $\pi$ we define: (i) the state transition probability matrix $\bm{P}^{\pi}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}=\mathbb{R}^{n\times n}$ where $\bm{P}^{\pi}(s,s^{+})=P\big{(}s^{+}|s,\pi(s)\big{)}$ for $s,s^{+}\in\mathcal{S}$, (ii) the state-action transition probability matrix $\overline{\bm{P}}^{\pi}\in\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|\times|\mathcal{S}\times\mathcal{A}|}=\mathbb{R}^{(nm)\times(nm)}$ where $\overline{\bm{P}}^{\pi}\big{(}(s,a),(s^{+},a^{+})\big{)}=P(s^{+}|s,a)$ if $a^{+}=\pi(s^{+})$ and $=0$ otherwise for $(s,a),(s^{+},a^{+})\in\mathcal{S}\times\mathcal{A}$, and (iii) the stage cost $\bm{c}^{\pi}\in\mathbb{R}^{|\mathcal{S}|}=\mathbb{R}^{n}$ where $\bm{c}^{\pi}(s)=\bm{c}\big{(}s,\pi(s)\big{)}$ for $s\in\mathcal{S}$.
Under policy $\pi$, the value function $\bm{v}^{\pi}\in\mathbb{R}^{|\mathcal{S}|}=\mathbb{R}^{n}$ is the expected discounted cost endured by following policy $\pi$ over an infinite-horizon trajectory, that is, for each $s\in\mathcal{S}$,
$\bm{v}^{\pi}(s)$ $\coloneqq\mathbb{E}_{s_{t+1}\sim\bm{P}^{\pi}(s_{t},\cdot)}\left[\sum_{t=0}^{\infty}\gamma^{t}\bm{c}^{\pi}(s_{t})\,\middle|\,s_{0}=s\right].$
The action-value function $\bm{q}^{\pi}\in\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|}=\mathbb{R}^{nm}$ (or the so-called Q-function) under policy $\pi$ is defined as follows: for each $(s,a)\in\mathcal{S}\times\mathcal{A}$
$\bm{q}^{\pi}(s,a)\coloneqq\bm{c}(s,a)+\gamma\ \mathbb{E}_{s^{+}\sim P(\cdot|s,a)}\left[\bm{v}^{\pi}(s^{+})\right].$
@@
The optimal value function can be characterized as the solution to the fixed-point equation $\bm{v}^{\star}=\mathbf{T}(\bm{v}^{\star})$, where $\mathbf{T}:\mathbb{R}^{|\mathcal{S}|}\rightarrow\mathbb{R}^{|\mathcal{S}|}$ is the so-called Bellman (optimality) operator defined as follows: for each $s\in\mathcal{S}$
$[\mathbf{T}(\bm{v}^{\star})](s)\coloneqq\min_{a\in\mathcal{A}}\left\{\bm{c}(s,a)+\gamma\ \mathds{E}_{s^{+}\sim P(\cdot|s,a)}\left[\bm{v}^{\star}(s^{+})\right]\right\}.$ (3)
@@
From the Banach fixed-point theorem (see, e.g., [27, Thm. 6.2.3]), the VI algorithm converges to $\bm{v}^{\star}$ with a linear rate $\gamma$. Correspondingly, one can derive the fixed-point characterization $\bm{q}^{\star}=\overline{\mathbf{T}}(\bm{q}^{\star})$ of the optimal Q-function, where
$[\overline{\mathbf{T}}(\bm{q}^{\star})](s,a)\coloneqq\bm{c}(s,a)+\gamma\ \mathds{E}_{s^{+}\sim P(\cdot|s,a)}\left[\min_{a^{+}\in\mathcal{A}}\bm{q}^{\star}(s^{+},a^{+})\right],$
for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ [31, Fact 3]. This characterization is particularly useful when one only has access to samples $\widehat{s}^{~{}+}\sim P(\cdot|s,a)$ of the next state $s^{+}$ for each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ (and not to the true transition probability kernel of the MDP). In particular, let us define the empirical Bellman operator $\widehat{\mathbf{T}}:\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|}\times\mathcal{S}\rightarrow\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|}$ as follows: for each $(s,a,s^{+})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$
$[\widehat{\mathbf{T}}(\bm{q},s^{+})](s,a)$ $\coloneqq\bm{c}(s,a)+\gamma\ \min_{a^{+}\in\mathcal{A}}\bm{q}(s^{+},a^{+}).$
@@
Above, $\lambda_{k}\geq 0$ are the step-sizes, The QL algorithm is also guaranteed to converge to $\bm{q}^{\star}$ almost surely given that the step-sizes satisfy the Robbins–Monro conditions (i.e., $\sum_{k=0}^{\infty}\lambda_{k}=\infty$ and $\sum_{k=0}^{\infty}\lambda^{2}_{k}<\infty$[32, 18]. In particular, with a polynomial step-size $\lambda_{k}=1/(1+k)^{\omega}$ with $\omega\in(\nicefrac{{1}}{{2}},1)$, QL outputs an $\epsilon$-accurate Q-function with high probability after $\tilde{\operatorname{\mathcal{O}}}(\tau^{-4/\omega}\cdot\epsilon^{-2/\omega}+\tau^{1/(1-\omega)})$ iterations of synchronous sampling with $\tau=1-\gamma$ [8], while with a re-scaled linear step-size $\lambda_{k}=1/(1+\tau k)$, QL has been shown to require $\tilde{\operatorname{\mathcal{O}}}(\tau^{-5}\cdot\epsilon^{-2})$ iterations of synchronous sampling for the same performance [33].
@@ 3. Rank-one value iteration (R1-VI)
$\pi^{\bm{v}}(s)\in\operatorname*{\arg\!\min}_{a\in\mathcal{A}}\mathds{E}_{s^{+}\sim P(\cdot|s,a)}\!\!\left[\bm{c}(s,a)+\gamma\bm{v}(s^{+})\right],\forall s\in\mathcal{S}.$ (6)
@@
is a rank-one approximation of the true transition probability matrix $\bm{P}_{k}$ at iteration $k$, with $\bm{d}_{k}\coloneqq\bm{d}_{\pi_{k}}\in\Delta(\mathcal{S})$ being a stationary distribution of the greedy policy $\pi_{k}=\pi_{\bm{v}_{k}}$, i.e., a solution of $\bm{d}_{k}^{\top}\bm{P}_{k}=\bm{d}_{k}^{\top}$. Under certain conditions, (10) is indeed “the best” rank-1 approximation of $\bm{P}_{k}$:
@@
$(\bm{I}-\gamma\widetilde{\bm{P}}_{k})^{-1}=(\bm{I}-\gamma\bm{1}\bm{d}_{k}^{\top})^{-1}=I+\tfrac{\gamma}{1-\gamma}\bm{1}\bm{d}_{k}^{\top},$
@@
$\bm{f}=\bm{P}_{k}^{\top}\bm{d}_{k}^{(i)},\ \bm{d}_{k}^{(i+1)}=\frac{\bm{f}}{\left\|\bm{f}\right\|_{1}},\ i=0,1,\ldots,I-1,$ (13)
@@
$(\bm{I}-\gamma\bm{P}_{k})^{-1}$ $=\!\sum_{\ell=0}^{\infty}\gamma^{\ell}\bm{P}_{k}^{\ell}\!=\sum_{\ell=0}^{L-1}\gamma^{\ell}\bm{P}_{k}^{\ell}+\gamma^{L}\bm{P}_{k}^{L}\sum_{\ell=0}^{\infty}\gamma^{\ell}\bm{P}_{k}^{\ell}$
$=\sum_{\ell=0}^{L-1}\gamma^{\ell}\bm{P}_{k}^{\ell}+\gamma^{L}\bm{P}_{k}^{L}(\bm{I}-\gamma\bm{P}_{k})^{-1}.$
@@
$\bm{G}_{k}$ $=\sum_{\ell=0}^{L-1}\gamma^{\ell}\bm{P}_{k}^{\ell}+\gamma^{L}\bm{P}_{k}^{L}(\bm{I}-\gamma\widetilde{\bm{P}}_{k})^{-1}$
$=\sum_{\ell=0}^{L}\gamma^{\ell}\bm{P}_{k}^{\ell}+\frac{\gamma^{L+1}}{1-\gamma}\bm{1}\bm{d}_{k}^{\top}.$
@@ 4. Rank-one Q-learning (R1-QL)
where $\overline{\bm{P}}_{k}\coloneqq\overline{\bm{P}}^{\pi^{\bm{q}_{k}}}$ is the state-action transition probability matrix of the MDP under the greed policy $\pi^{\bm{q}_{k}}$.
@@
$a^{+}=\widehat{a}_{k}^{~{}+}=\operatorname*{\arg\!\min}\limits_{a^{+}\in\mathcal{A}}\bm{q}_{k}(\widehat{s}^{~{}+}_{k},a^{+}).$
@@
Now, using the approximation $\widehat{\bm{d}}_{k-1}\approx\widehat{\bm{P}}_{k-1}^{\top}\widehat{\bm{d}}_{k-1}$ (i.e., assuming $\widehat{\bm{d}}_{k-1}$ is the stationary distribution of $\widehat{\bm{P}}_{k-1}$ which does not hold exactly since $\widehat{\bm{d}}_{k-1}$ is only an approximation of the stationary distribution of $\widehat{\bm{P}}_{k-1}$), we derive,
$\widehat{\bm{d}}_{k}=(1-\lambda_{k})\widehat{\bm{d}}_{k-1}+\lambda_{k}\bm{F}_{k}^{\top}\widehat{\bm{d}}_{k-1}$
or, equivalently, for each $(s,a)\in\mathcal{S}\times\mathcal{A}$
$\widehat{\bm{d}}_{k}(s,a)=(1-\lambda_{k})\widehat{\bm{d}}_{k-1}(s,a)+\lambda_{k}\widehat{\bm{d}}_{k-1}(s^{+},a^{+}),$ (18)
@@
8:        $\bm{f}(s,a)=(1-\lambda_{k})\widehat{\bm{d}}_{k-1}(s,a)+\lambda_{k}\widehat{\bm{d}}_{k-1}(s^{+},a^{+})$;

### v2

@@ Abstract.
1Delft Center for Systems and Control, Delft University of Technology, The Netherlands. 2Department of Mechanical and Industrial Engineering, University of Toronto, Canada. Correspondence to: Arman S. Kolarijani $<$a.sharifikolarijani@tudelft.nl$>$.
This work was partially supported by the Horizon Europe Pathfinder Open project RELIEVE-101099481 and by the European Research Council (ERC) project TRUST-949796.
The authors would like to thank the reviewers of ICML 2025 for their useful comments.
@@ 1. Introduction
Contribution. In this paper, we propose a novel algorithm that modifies the VI algorithm by incorporating a computationally efficient PI-type update rule. To this end, we consider the update rule (1) with a matrix gain of the form $\bm{G}_{k}=(\bm{I}-\widetilde{\bm{P}}_{k})^{-1}$, where $\widetilde{\bm{P}}_{k}$ is a rank-one approximation of $\bm{P}_{k}$. To be precise, we consider the approximation $\widetilde{\bm{P}}_{k}=\bm{1}\bm{d}_{k}^{\top}$, where $\bm{d}_{k}=\bm{P}_{k}^{\top}\bm{d}_{k}$ is a stationary distribution of $\bm{P}_{k}$ and $\bm{1}$ is the all-one vector. The proposed algorithm then uses the power method for approximating $\bm{d}_{k}$ iteratively using the true matrix $\bm{P}_{k}$ in the planning problem and its sampled version in the learning problem. In particular,
@@
    we compare the proposed R1-VI and R1-QL algorithms with the state-of-the-art algorithms for solving planning and learning problems of MDPs and show the empirically faster convergence of the proposed algorithms compared to the ones with the same per-iteration computational complexity (i.e., the first-order algorithms and their accelerated versions).
@@ 2. Optimal control of MDPs
Consider a finite MDP $(\mathcal{S},\mathcal{A},P,c,\gamma)$. Here, $\mathcal{S}:=\{1,2,\ldots,n\}$ and $\mathcal{A}:=\{1,2,\ldots,m\}$ are the state and action spaces, respectively. The transition kernel $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the conditional probability $P(s^{+}|s,a)$ of the transition to state $s^{+}$ given the current state-action pair $(s,a)$. The function $\bm{c}\in\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|}=\mathbb{R}^{nm}$, bounded from below, represents the stage cost $\bm{c}(s,a)$ of taking the control action $a$ while the system is in state $s$. And, $\gamma\in(0,1)$ is the discount factor which can be seen as a trade-off parameter between short- and long-term costs.
A control policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ is a mapping from states to actions. Fix policy $\pi$. For the corresponding Markov chain under the policy $\pi$ we define:
  • (i)
    the state transition probability matrix $\bm{P}^{\pi}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}=\mathbb{R}^{n\times n}$ where $\bm{P}^{\pi}(s,s^{+})=P\big{(}s^{+}|s,\pi(s)\big{)}$ for $s,s^{+}\in\mathcal{S}$;
  • (ii)
    the state-action transition probability matrix $\overline{\bm{P}}^{\pi}\in\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|\times|\mathcal{S}\times\mathcal{A}|}=\mathbb{R}^{(nm)\times(nm)}$ where
    $\overline{\bm{P}}^{\pi}\big{(}(s,a),(s^{+},a^{+})\big{)}=\left\{\begin{array}[]{ll}P(s^{+}|s,a)&\text{if }a^{+}=\pi(s^{+}),\\ 0&\text{otherwise},\end{array}\right.\quad\text{for }(s,a),(s^{+},a^{+})\in\mathcal{S}\times\mathcal{A}.$
  • (iii)
    the stage cost $\bm{c}^{\pi}\in\mathbb{R}^{|\mathcal{S}|}=\mathbb{R}^{n}$ where $\bm{c}^{\pi}(s)=\bm{c}\big{(}s,\pi(s)\big{)}$ for $s\in\mathcal{S}$.
Under policy $\pi$, the value function $\bm{v}^{\pi}\in\mathbb{R}^{|\mathcal{S}|}=\mathbb{R}^{n}$ is the expected discounted cost endured by following policy $\pi$ over an infinite-horizon trajectory, that is,
$\bm{v}^{\pi}(s)$ $\coloneqq\mathbb{E}_{s_{t+1}\sim\bm{P}^{\pi}(s_{t},\cdot)}\left[\sum_{t=0}^{\infty}\gamma^{t}\bm{c}^{\pi}(s_{t})\,\middle|\,s_{0}=s\right],\quad\forall s\in\mathcal{S}.$
The action-value function $\bm{q}^{\pi}\in\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|}=\mathbb{R}^{nm}$ (or the so-called Q-function) under policy $\pi$ is defined as
$\bm{q}^{\pi}(s,a)\coloneqq\bm{c}(s,a)+\gamma\ \mathbb{E}_{s^{+}\sim P(\cdot|s,a)}\left[\bm{v}^{\pi}(s^{+})\right],\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$
@@
The optimal value function can be characterized as the solution to the fixed-point equation $\bm{v}^{\star}=\mathbf{T}(\bm{v}^{\star})$, where $\mathbf{T}:\mathbb{R}^{|\mathcal{S}|}\rightarrow\mathbb{R}^{|\mathcal{S}|}$ is the so-called Bellman (optimality) operator defined as follows:
$[\mathbf{T}(\bm{v}^{\star})](s)\coloneqq\min_{a\in\mathcal{A}}\left\{\bm{c}(s,a)+\gamma\ \mathds{E}_{s^{+}\sim P(\cdot|s,a)}\left[\bm{v}^{\star}(s^{+})\right]\right\},\quad\forall s\in\mathcal{S}.$ (3)
@@
From the Banach fixed-point theorem (see, e.g., [27, Thm. 6.2.3]), the VI algorithm converges to $\bm{v}^{\star}$ with a linear rate $\gamma$. Correspondingly, one can derive the fixed-point characterization $\bm{q}^{\star}=\overline{\mathbf{T}}(\bm{q}^{\star})$ of the optimal Q-function, where [31, Fact 3]
$[\overline{\mathbf{T}}(\bm{q}^{\star})](s,a)\coloneqq\bm{c}(s,a)+\gamma\ \mathds{E}_{s^{+}\sim P(\cdot|s,a)}\left[\min_{a^{+}\in\mathcal{A}}\bm{q}^{\star}(s^{+},a^{+})\right],\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$
This characterization is particularly useful when one only has access to samples $\widehat{s}^{~{}+}\sim P(\cdot|s,a)$ of the next state $s^{+}$ for each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ (and not to the true transition probability kernel of the MDP). In particular, let us define the empirical Bellman operator $\widehat{\mathbf{T}}:\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|}\times\mathcal{S}\rightarrow\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|}$ as follows:
$[\widehat{\mathbf{T}}(\bm{q},s^{+})](s,a)$ $\coloneqq\bm{c}(s,a)+\gamma\ \min_{a^{+}\in\mathcal{A}}\bm{q}(s^{+},a^{+}),\quad\forall(s,a,s^{+})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}.$
@@
Above, $\lambda_{k}\geq 0$ are the step-sizes. The QL algorithm is also guaranteed to converge to $\bm{q}^{\star}$ almost surely given that the step-sizes satisfy the Robbins–Monro conditions (i.e., $\sum_{k=0}^{\infty}\lambda_{k}=\infty$ and $\sum_{k=0}^{\infty}\lambda^{2}_{k}<\infty$[32, 18]. In particular, with a polynomial step-size $\lambda_{k}=1/(1+k)^{\omega}$ with $\omega\in(\nicefrac{{1}}{{2}},1)$, QL outputs an $\epsilon$-accurate Q-function with high probability after $\tilde{\operatorname{\mathcal{O}}}(\tau^{-4/\omega}\cdot\epsilon^{-2/\omega}+\tau^{1/(1-\omega)})$ iterations of synchronous sampling with $\tau=1-\gamma$ [8], while with a re-scaled linear step-size $\lambda_{k}=1/(1+\tau k)$, QL has been shown to require $\tilde{\operatorname{\mathcal{O}}}(\tau^{-5}\cdot\epsilon^{-2})$ iterations of synchronous sampling for the same performance [33].
@@ 3. Rank-one value iteration (R1-VI)
$\pi^{\bm{v}}(s)\in\operatorname*{\arg\!\min}_{a\in\mathcal{A}}\mathds{E}_{s^{+}\sim P(\cdot|s,a)}\!\!\left[\bm{c}(s,a)+\gamma\bm{v}(s^{+})\right],\quad\forall s\in\mathcal{S}.$ (6)
@@
is a rank-one approximation of the true transition probability matrix $\bm{P}_{k}$ at iteration $k$, with $\bm{d}_{k}\coloneqq\bm{d}_{\pi_{k}}\in\Delta(\mathcal{S})$ being a stationary distribution of the greedy policy $\pi_{k}=\pi^{\bm{v}_{k}}$, i.e., a solution of $\bm{d}_{k}^{\top}\bm{P}_{k}=\bm{d}_{k}^{\top}$. Under certain conditions, (10) is indeed “the best” rank-1 approximation of $\bm{P}_{k}$:
@@
$(\bm{I}-\gamma\widetilde{\bm{P}}_{k})^{-1}=(\bm{I}-\gamma\bm{1}\bm{d}_{k}^{\top})^{-1}=\bm{I}+\tfrac{\gamma}{1-\gamma}\bm{1}\bm{d}_{k}^{\top},$
@@
$\bm{f}=\bm{P}_{k}^{\top}\bm{d}_{k}^{(i)},\ \bm{d}_{k}^{(i+1)}=\frac{\bm{f}}{\left\|\bm{f}\right\|_{1}},\quad i=0,1,\ldots,I-1,$ (13)
@@
$(\bm{I}-\gamma\bm{P}_{k})^{-1}=\!\sum_{\ell=0}^{\infty}\gamma^{\ell}\bm{P}_{k}^{\ell}\!=\sum_{\ell=0}^{L-1}\gamma^{\ell}\bm{P}_{k}^{\ell}+\gamma^{L}\bm{P}_{k}^{L}\sum_{\ell=0}^{\infty}\gamma^{\ell}\bm{P}_{k}^{\ell}=\sum_{\ell=0}^{L-1}\gamma^{\ell}\bm{P}_{k}^{\ell}+\gamma^{L}\bm{P}_{k}^{L}(\bm{I}-\gamma\bm{P}_{k})^{-1}.$
@@
$\bm{G}_{k}=\sum_{\ell=0}^{L-1}\gamma^{\ell}\bm{P}_{k}^{\ell}+\gamma^{L}\bm{P}_{k}^{L}(\bm{I}-\gamma\widetilde{\bm{P}}_{k})^{-1}=\sum_{\ell=0}^{L}\gamma^{\ell}\bm{P}_{k}^{\ell}+\frac{\gamma^{L+1}}{1-\gamma}\bm{1}\bm{d}_{k}^{\top}.$
@@ 4. Rank-one Q-learning (R1-QL)
where $\overline{\bm{P}}_{k}\coloneqq\overline{\bm{P}}^{\pi^{\bm{q}_{k}}}$ is the state-action transition probability matrix of the MDP under the greed policy $\pi^{\bm{q}_{k}}(s)\in\operatorname*{\arg\!\min}_{a\in\mathcal{A}}\bm{q}_{k}(s,a)$ for $s\in\mathcal{S}$.
@@
$a^{+}=\widehat{a}_{k}^{~{}+}\in\operatorname*{\arg\!\min}\limits_{a^{+}\in\mathcal{A}}\bm{q}_{k}(\widehat{s}^{~{}+}_{k},a^{+}).$
@@
Now, using the approximation $\widehat{\bm{d}}_{k-1}\approx\widehat{\bm{P}}_{k-1}^{\top}\widehat{\bm{d}}_{k-1}$ (i.e., assuming $\widehat{\bm{d}}_{k-1}$ is a stationary distribution of $\widehat{\bm{P}}_{k-1}$ which does not hold exactly since $\widehat{\bm{d}}_{k-1}$ is only an approximation of a stationary distribution of $\widehat{\bm{P}}_{k-1}$), we derive,
$\widehat{\bm{d}}_{k}=(1-\lambda_{k})\widehat{\bm{d}}_{k-1}+\lambda_{k}\bm{F}_{k}^{\top}\widehat{\bm{d}}_{k-1},$
or, equivalently,
$\widehat{\bm{d}}_{k}(s,a)=(1-\lambda_{k})\widehat{\bm{d}}_{k-1}(s,a)+\lambda_{k}\widehat{\bm{d}}_{k-1}(s^{+},a^{+}),\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A},$ (18)
@@
8:        $\bm{f}(s,a)=(1-\lambda_{k})\widehat{\bm{d}}_{k-1}(s,a)+\lambda_{k}\widehat{\bm{d}}_{k-1}(\widehat{s}^{~{}+}_{k},\widehat{a}_{k}^{~{}+})$;